OpenStudy (anonymous):

John is having computer problems. He noticed that 6 of the 26 letters on his keyboard aren’t working. If John’s computer password consists of 5 distinct letters, what is the probability that John will still be able to type in his password (all 5 letters are working)?

OpenStudy (shane_b):

I'm going to take a stab at this one...but I'm no probability expert so someone jump in if you seem something wrong. You have to multiply the odds of each character working... 1st char: 20/26 2nd char: 19/26 3rd char: 18/26 4th char: 17/26 5th char: 16/26 \[\frac{20}{26}*\frac{19}{26}*\frac{18}{26}*\frac{17}{26}*\frac{16}{26}=15.7\%\]

OpenStudy (anonymous):

What the answer key says is that the answer is 3876/16445

OpenStudy (shane_b):

3876/16445 is a 23.6% chance...I don't see where my mistake is but maybe someone else will.

OpenStudy (anonymous):

Well, here's how they got it. :/

OpenStudy (anonymous):

I guess it makes sense now huh?

OpenStudy (shane_b):

Ah yes...silly mistake thinking that the 26 would remain constant in each calculation!

OpenStudy (shane_b):

I should have realized that he would obviously not try to use the broken keys

OpenStudy (shane_b):

If he was blind, then I'm certain my answer would be right :P

OpenStudy (anonymous):

Yeah, I waited to look at the solution methods because I like trying to figure it out somewhat. I'm terrible at probability, but I should've figured it out :) lol. THanks for trying though!

OpenStudy (shane_b):

I learned something so it wasn't wasted time :)